OpenStudy (anonymous):

Hello everyone! Solve for x: -3|x + 7| = -12 A)x = 5 over 3, x = −19 over 3 B)x = −3, x = −11 C)x = −3, x = 11 D)No solution

OpenStudy (anonymous):

start by dividing both sides by \(-3\) to get \[|x+7|=4\] then solve two equations \[x+7=4\] and also solve \[x+7=-4\]

OpenStudy (godlovesme):

is this the one u need help with? @extremearianna

OpenStudy (anonymous):

yes @Godlovesme

OpenStudy (godlovesme):

k wait a min :)

OpenStudy (anonymous):

@satellite73 thank you. I solved the other two equations and I got -3 & -11! So the correct answer is B :-D

OpenStudy (godlovesme):